OpenStudy (ams98):

Tony is trying to find the equation of a quadratic that has a focus of (−1, 4) and a directrix of y = 8. Describe to Tony your preferred method for deriving the equation. Make sure you use Tony's situation as a model to help him understand.

OpenStudy (aum):

A parabola has the property that any point on the parabola is equidistant from the focus and the directrix. Sp to find the equation of the parabola (which is a quadratic) take a general point (x,y) on the parabola, find its distance from the the focus, find its distance from the directrix, equate them and simplify.

OpenStudy (ams98):

But how do I do that with this problem specifically.

OpenStudy (aum):

Equating the distance is same as equating the square of the distance. The square of the distance between a general point on the parabola, (x,y) and the focus, (-1, 4) is: \((x - -1)^2 + (y-4)^2 = (x+1)^2 + (y-4)^2\) The square of the distance between (x,y) and the directrix y = 8 is: \((y - 8)^2\) Equate the square of the distances: \((x+1)^2 + (y-4)^2 = (y-8)^2\) Simplify.

OpenStudy (aum):

\((x+1)^2 + (y-4)^2 = (y-8)^2\) \((x+1)^2 = (y-8)^2 - (y-4)^2 \) \((x+1)^2 = (y-8+y-4)(y-8-y+4) \) \((x+1)^2 = (2y-12)(-4) = 2(y-6)(-4) = -8(y-6)\) \(\large \frac{-1}{8}(x+1)^2 = (y-6) \) \(\large \frac{-1}{8}(x+1)^2 + 6 = y \) \(\large y = \frac{-1}{8}(x+1)^2 + 6\)

OpenStudy (ams98):

holy crap is the anwser ? I was doing it SO wrong

OpenStudy (aum):

I have to double-check...

OpenStudy (ams98):

Thank you so much! Wow

OpenStudy (aum):

Yes. \(\large y = -\frac{1}{8}(x+1)^2 + 6\) is the correct answer.